Model Description

We use the NEMO (Nucleus for European Modelling of the Ocean), an OGCM based on primitive equations [Madec et al., 1999; Madec, 2008].

Ocean model:

NEMO is an ocean modelling framework which is composed of three ”engines.” The ”engines” provide numerical solutions of ocean, sea-ice, tracers and biochemistry equations and their related physics (http://www.nemo-ocean.eu/About-NEMO). It is intended to be a flexible tool for studying the ocean and its interactions with the others components of the earth climate system such as atmosphere, sea-ice, and biogeochemical tracers over a wide range of space and time scales.

We intend to use the ORCA2_{L}IM_{P} ISCESconfigurationofNEMOconsistingof3majorengines, theoceanmodel, ORCA2, coupledtoanicemodel, LIM, andabiogeochemichalmodel, P ISCES.

ORCA2 is a primitive equation model adapted to regional and global ocean circulation problems. Prognostic variables are the three-dimensional velocity field, the sea surface height, the temperature and the salinity.

In the ORCA2_{L}IM_{P} ISCESconfiguration, themodeldomainextendsfrom78Sto90N.T hemodelusesaglobaltripolarorthogonalcurvilineargridwith2degzonalresolution, andthemeridionalresolutionisenhancedto0.5degattheequatorandathighlatitudesto2cos(lat)degsouthof20S.T hehorizontalgridfeaturestwopointsofconvergenceintheNorthernHemisphere, bothsituatedoncontinentstoovercometheNorthP olesingularityfoundforgeographicalmeshes, thusavoidssingularitypointinsidethecomputationaldomain[MadecandImbard, 1996], andtheratioofanisotropyisnearlyoneeverywhere.Minimumresolutioninhighlatitudesisabout65kmintheArcticand50kmintheAntarctic[T immermannetal, 2005]. Other configurations with a higher resolution (1 deg, 0.5 deg, and 0.25 deg) have also been developed

[DRAKKAR group, 2007].

In the vertical direction, the model uses a full or partial step z-coordinate, or s-coordinate, or a mixture of the two. There are 31 levels, with 10 levels in the top 100m and partial steps in the lowest level. The vertical mesh is deduced from a mathematical function of z [Madec and Imbard 1996].

The distribution of variables is a three-dimensional Arakawa C-type grid.

Various physical choices are available to describe ocean physics, including TKE and KPP vertical physics. Lateral tracer mixing is done along isopycnals. Eddy induced tracer advection is parameterized following

Gent and McWilliams (1990).

Horizontal momentum is mixed along model level surfaces using the eddy viscosity coefficients varying with latitude, longitude and depth. The diffusion on tracers acts along the isopycnal surfaces (neutral surface).

Vertical eddy diffusivity and viscosity coefficients are calculated using a 1.5 order turbulent kinetic energy model [Gaspar et al., 1990].

Zero fluxes of heat and salt and no-slip conditions are applied through lateral solid boundaries. At the bottom boundary, zero fluxes of heat and salt are applied through the ocean bottom.

The ocean is subject to the external forcings of heat, freshwater, and momentum fluxes from the atmosphere and/or the sea-ice. The solar radiation penetrates the top meters of the ocean. The downward irradiance I(z) is formulated with two extinction coefficients [Paulson and Simpson, 1977], whose values correspond to a Type I water, the most transparent water, in Jerlov’s classification [Jerlov, 1968]. We intend to adopt in the model the parameterization proposed by Morel and Antoine [1994] to improve the upper ocean solar heating rate affected by the abundance of phytoplankton [Nakamoto, et al, 2001; Ueyoshi, et al, 2003; Ueyoshi, et al, 2005; Frouin, et al, 2007; Subrahmanyam, et al, 2007].

Solar radiation absorption and local heating within the upper layers of the ocean are strongly influenced by the abundance of phytoplankton. Depending on the degree of the chlorophyll concentration, the heat deposition occurs within a layer that may vary in thickness from less than 10 m to more than 100 m. The parameterization allows the vertical profiles of heating rate to be predicted from the phytoplanktonic pigment concentration. For the infrared waveband, not influenced by biological materials, the irradiance profile is described by a single exponential function. For the ultraviolet and visible (¡ 750 nm) band, a bimodal exponential form is adopted. The weights associated with each of these exponential functions, along with their specific attenuation lenghts, are dependent upon pigment concentration. These dependencies are explicated through polynomial formulas. Pigment values predicted by the PISCES model are used for the heat rate calculation (Morel and Antoine, 1994).

Sea-ice model:

Within NEMO, the ocean is interfaced with the interactive sea-ice model LIM (Louvain-la-Neuve sea ice model) [Fichefet and Morales-Maqueda, 1997, Timmermann, et al, 2005]. LIM3, the latest version of LIM, is a C-grid dynamicthermodynamic model, including the representation of the subgrid-scale distributions of five-category sea ice thickness, enthalpy, salinity and age. Brine entrapment and drainage as well as brine impact on ice thermodynamics are explicitly included (Vancoppenolle, et al, 2009). LIM3 is found to